The waves on a string have wave speed 8.00m/s, amplitude 0.0600m, and wavelength 

0.400m. The waves travel in the -x direction and at t = 0s, the end of the string has its maximum 

upward displacement. 

(a) Write a wave function describing the wave and

(b). Find the transverse displacement of a particle at x = 0.350m at time t = 0.150s


1
Expert's answer
2022-04-11T12:08:30-0400

(a) Write a wave function describing the wave

"y(x,t)=A\cos(2\pi ft+\frac{2\pi}{\lambda}x)"

"f=v/\lambda=8.0/0.400=20\:\rm Hz"

"y(x,t)=0.0600{\:\rm m}\cos(40\pi t+5\pi x)"

(b) Find the transverse displacement of a particle at x = 0.350m at time t = 0.150s

"y=0.0600{\:\rm m}\cos(40\pi *0.150+5\pi *0.350)\\
=0.0424\:\rm m"